# Unconventional criticality in $O(D)$-invariant loop-constrained Landau theory

### v1

@@ Abstract
We study the critical behavior of a Landau-type theory for ferroelectrics in which the polarization order parameter $\bm{P}$ is subject to a divergence-free constraint, such that only loop-like polarization configurations contribute to the partition function. This constraint forces the number of components of the field $\bm{P}$ to equal the spatial dimensionality $D$, allowing a natural extension of the present theory to $O(N)$ models with $N=D$. Renormalization group analysis reveals critical behavior beyond the conventional Landau-Ginzburg-Wilson paradigm. In particular, the order parameter acquires a remarkably large anomalous dimension, with $\eta\approx 0.239$ in three dimensions – significantly exceeding the value $\eta\approx 0.034$ typical of the $O(3)$ universality class. This is due to an emergent gauge symmetry originating with the local constraint.
@@
Introduction – The Landau paradigm of phase transitions, built upon the concept of spontaneous symmetry breaking and local order parameters, has been a cornerstone of condensed matter physics. However, numerous systems exhibiting topological order, emergent gauge fields, and fractionalized excitations have revealed the limitations of this framework [1, 2, 3, 4], triggering the exploration of critical phenomena that go beyond conventional symmetry-breaking descriptions. These include transitions driven by topological order, quantum entanglement, and fractionalized excitations, as seen in quantum spin liquids and deconfined quantum critical points. Such transitions challenge our understanding of universality and scaling, offering insights into complex phases of matter with potential applications in quantum computation, quantum information, and quantum sensing, as well as in materials science.
@@
In particular, systems governed by local constraints offer a promising platform for unconventional criticality. A prominent example involves divergence-free conditions on vector fields, which manifest across various physical systems. For instance, in frustrated antiferromagnets on pyrochlore and kagome lattices [5], divergence-free constraints on staggered magnetization or local flux variables emerge from local constraints imposed by competing interactions. Such a constraint also arises in the dual theories of three-dimensional quantum dimer models [6, 7].
@@
In continuum systems, magnetostatics offers a classical example, where the solenoidal constraint on magnetic field, ${\mbox{$\nabla$}}\cdot\bm{B}=0$, governs flux conservation and enforces loop-like field configurations [8]. This imposes topological restrictions on the field configurations in magnetic fluids, leading to conserved magnetic helicity and the formation of coherent, filamentary magnetic structures [9]. In the context of incompressible fluids, the velocity field $\bm{v}$ satisfies ${\mbox{$\nabla$}}\cdot\bm{v}=0$, leading to nonlocal interactions through pressure mediation. According to Arnold’s theorem for incompressible liquids [10], the stationary fluid flow in this case is composed of two types of elementary topological excitations: either vortex-like or fibered into tori – Hopfions. Similarly, in nanoscale ferroelectrics, e.g. in SrTiO3/PbTiO3 superlattices or PbTiO3 and PbZr0.6Ti0.4O3 nanoparticles [11, 12, 13, 14], a zero-divergence constraint on the polarization field ${\mbox{$\nabla$}}\cdot\bm{P}=0$ plays a central role in the formation of vortex-like and Hopfion-like textures, in which the polarization field streamlines wind around nested cylinders or tori. This happens to avoid the formation of polarization bound charges, $\rho_{p}=-{\mbox{$\nabla$}}\cdot\bm{P}$, arising due to the nonuniform distribution of polarization dipoles, and to minimize the huge electrostatic energy associated with depolarization effects. Such constraint-driven structuring of field lines gives rise to rich topological phenomena [15, 16, 17, 14], with analogies to hydrodynamics and emergent gauge theories. In superfluids, the constraint ${\mbox{$\nabla$}}\cdot\bm{v}_{s}=0$ on the superfluid velocity $\bm{v}_{s}$ ultimately leads to vortex loop excitations [18, 19, 20], which can be used to analyze the phase structure [21, 22]. In two dimensions, where vortices are points, this point of view leads to the celebrated Berezinskii-Kosterlitz-Thouless (BKT) phase transition [23, 24].
@@
These are just a few examples of the diverse classical and quantum systems where divergence-free conditions govern the low-energy physics, generating gauge structures and long-range correlations. When supplemented with continuous symmetries such as $O(D)$, these systems can exhibit novel, Landau-incompatible critical behavior, which is in the focus of this Letter. As an exemplary system to reveal the constraint-imposed effects, we consider ferroelectric materials with zero divergence of polarization, ${\mbox{$\nabla$}}\cdot\bm{P}=0$, giving rise to topological configurations that influence the system’s critical properties. Further, we refer to such materials as topological ferroelectrics.
@@
A natural question that emerges from the above considerations concerns the role played by the zero divergence constraint in determining the phase structure of the different physical systems. How much does this constraint affect the Landau-Ginzburg-Wilson (LGW) theory of a given system, like, for instance, a ferroelectric? Remarkably, it turns out that the critical behavior of correlation functions is affected significantly, reflecting a scenario that is usually encountered in physical systems exhibiting fractionalization. This is a property underlying the decay of an order parameter into emergent building blocks, similarly to the physics of deconfinement as studied in high-energy physics. In condensed matter physics, it is typically associated to an emerging gauge symmetry, with realizations in quantum antiferromagnets featuring competing orders. The outcome is a phase structure that is incompatible with the LGW paradigm. One of the most salient features of the so called deconfined quantum critical points (DQCs) is a unusually large anomalous dimension, $\eta$, of the order parameter $O(\bm{x})$ [1]. In LGW-compatible systems $\eta$ is typically much smaller than the thermodynamic critical exponents, namely, those associated to the critical singularities of the free energy density. The large anomalous dimension indicates that the order parameter is actually made of more elementary constituents, like for example spinons in quantum antiferromagnets. In this Letter, we show that a LGW theory accounting for the constraint ${\mbox{$\nabla$}}\cdot\bm{P}=0$ also leads to a large anomalous dimension, with no obvious presence of fractionalized excitations. In fact, we find at two-loop order in perturbation theory that $\eta=0.239$ in three dimensions. Recall that in an LGW theory with $O(N)$ symmetry the one-loop result vanishes, so it is necessary to go beyond one-loop order to find a non-vanishing value of $\eta$, which is typically one order of magnitude smaller than the value we find here. For example, in a $O(3)$ symmetric theory, which is the natural case to compare with in the present context, the result is $\eta\approx 0.034$. Crucially, the constraint ${\mbox{$\nabla$}}\cdot\bm{P}=0$ naturally imposes that $N=D$.
A structurally similar situation occurs in the nonlinear $O(N)$ $\sigma$-model, where the field $\bm{n}$ is constrained to a fixed-radius manifold, $\bm{n}^{2}=1$. This constraint also leads to significantly different critical exponents as compared to those for $O(N)$ universality class as given by the LGW paradigm. The similar consequences of both constraints may indicate the presence of additional divergence-free condition in $\sigma$-models, that are not manifest in the standard continuum formulation.
In our case, the zero divergence constraint can be understood as a local conservation law corresponding to a local gauge symmetry, since ${\mbox{$\nabla$}}\cdot\bm{P}=0$ can be solved as $\bm{P}={\mbox{$\nabla$}}\times\bm{A}$. This leads to the emergence of a gauge field and therefore to a large value of $\eta$. Usually, the argument runs the opposite way, with a gauge symmetry yielding a local conservation law.
@@
Substituting the fixed point values, $\hat{a}_{*}=-25/77$ and $\hat{b}_{*}=5/17$, calculated from Eqs. (9) for $D=3$, we obtain $\eta=0.239$. This value is significantly larger than the ones typically obtained from the LGW theory, thus highlighting the impact of the constraint ${\mbox{$\nabla$}}\cdot\bm{P}=0$.
@@
It remains to be seen whether the universality class uncovered here can be connected to some fractionalization mechanism, by representing the polarization $\bm{P}$ as a composite operator, similarly to the CP1 representation in quantum antiferromagnets [1]. Deconfined quantum criticality is indeed facilitated by a mechanism where staggered Berry phases interfere with quadrupled instanton events, which ultimately causes the order parameter to fall apart into spinons [1, 4].
Conclusions and final remarks – In this Letter we have studied the emergence of a LGW-incompatible universality class following from the effective action for a ferroelectric system with the additional constraint ${\mbox{$\nabla$}}\cdot\bm{P}=0$. This leads to a theory describing a partition function corresponding to an ensemble of polarization loops. Although the bare action itself resembles a standard Ginzburg-Landau theory, the constraint in the path integral measure significantly affects the critical behavior of the system. This leads to a feature usually encountered in theories for DQC [1], namely, a large anomalous dimension for the order parameter correlator. Remarkably, in this case there is no obvious fractionalization of the order parameter. Nevertheless, an emergent $U(1)$ gauge symmetry – another feature occurring in DQC – also arises here, since the constraint ${\mbox{$\nabla$}}\cdot\bm{P}=0$ can be solved as $\bm{P}={\mbox{$\nabla$}}\times\bm{A}$, turning the action manifestly gauge invariant. The resulting theory is reminiscent of an effective gauge theory for three-dimensional quantum dimer models [6, 7].

### v2

@@ Abstract
We study an unconventional phase transition in ferroelectrics where the polarization field is constrained to be divergence-free, allowing only loop-like configurations. This local constraint fundamentally alters the critical behavior, driving the system beyond the Landau-Ginzburg-Wilson paradigm. A renormalization group analysis shows that the polarization acquires an unusually large anomalous dimension, $\eta\approx 0.239$ in three dimensions, far exceeding the typical values in $O(3)$-invariant systems. We attribute this effect to an emergent gauge symmetry originating from the zero divergence constraint. Such gauge-field behavior is reminiscent of fractionalized phases, revealing a fundamental connection between constrained ferroelectrics and emergent gauge phenomena in correlated matter.
@@
Introduction – The Landau-Ginzburg-Wilson (LGW) paradigm of phase transitions, built upon the concept of spontaneous symmetry breaking and local order parameters, has been a cornerstone of condensed matter physics. However, numerous systems have revealed the limitations of this framework, triggering the exploration of critical phenomena that go beyond conventional symmetry-breaking descriptions. These include transitions driven by topological order, quantum entanglement, and fractionalized excitations, as seen in quantum spin liquids and deconfined quantum criticality (DQC) [1, 2, 3, 4]. Such transitions challenge our understanding of universality and scaling, offering insights into complex phases of matter with potential applications in quantum computation, quantum information, and quantum sensing, as well as in materials science.
@@
In particular, systems subjected to local constraints offer a promising platform for unconventional criticality. A prominent example involves divergence-free conditions on vector fields, which manifest across various physical systems. For instance, in frustrated antiferromagnets on pyrochlore and kagome lattices [5], divergence-free constraints on staggered magnetization or local flux variables emerge from local constraints imposed by competing interactions. Such a constraint also arises in theories of three-dimensional quantum dimer models [6, 7].
@@
In continuum systems, magnetostatics offers a classical example, where the solenoidal constraint on magnetic field, ${\mbox{$\nabla$}}\cdot\bm{B}=0$, governs flux conservation and enforces loop-like field configurations [8]. This imposes topological restrictions on magnetic fluids, leading to conserved magnetic helicity and the formation of coherent, filamentary magnetic structures [9]. In the context of incompressible fluids, the velocity field $\bm{v}$ satisfies ${\mbox{$\nabla$}}\cdot\bm{v}=0$, leading to nonlocal interactions through pressure mediation. According to Arnold’s theorem for incompressible liquids [10], the stationary fluid flow in this case is composed of two types of elementary topological excitations: either vortex-like or knotted hopfion-like ones. Similarly, in nanoscale ferroelectrics, e.g. in SrTiO3/PbTiO3 superlattices or PbTiO3 and PbZr0.6Ti0.4O3 nanoparticles [11, 12, 13, 14], a zero divergence constraint on the polarization field ${\mbox{$\nabla$}}\cdot\bm{P}=0$ plays a central role in the formation of vortex- and hopfion-like textures, in which the polarization field streamlines wind around nested cylinders or tori. This prevents the formation of polarization bound charges, $\rho_{p}=-{\mbox{$\nabla$}}\cdot\bm{P}$, which arise due to the nonuniform distribution of polarization dipoles, and minimizes the huge electrostatic energy associated with depolarization effects. Such constraint-driven structuring of field lines gives rise to rich topological phenomena [15, 16, 17, 14], with analogies to hydrodynamics and emergent gauge theories. Similarly, in a superfluid, the constraint ${\mbox{$\nabla$}}\cdot\bm{v}_{s}=0$ on the superfluid velocity $\bm{v}_{s}$ ultimately leads to vortex loop excitations [18, 19, 20], which determine its phase structure [21, 22]. In particular, in two dimensions, where vortices are points, this point of view leads to the celebrated Berezinskii-Kosterlitz-Thouless (BKT) phase transition [23, 24].
@@
These are just a few examples of the diverse classical and quantum systems where divergence-free conditions govern the low-energy physics, generating gauge structures and long-range correlations. When supplemented with continuous symmetries such as the orthogonal group $O(D)$, where $D$ is the spatial dimension, these systems can exhibit LGW-incompatible critical behavior, which is in the focus of this Letter. As an exemplary system revealing the constraint-imposed effects, we consider ferroelectric materials where the polarization is divergence-free, ${\mbox{$\nabla$}}\cdot\bm{P}=0$, giving rise to topological configurations that influence the system’s critical properties. We refer to such materials as topological ferroelectrics.
@@
A natural question emerging from the above considerations concerns the role played by the zero divergence constraint in determining the phase structure of the physical system. How much does such a constraint affect the LGW theory of a ferroelectric system? Remarkably, we find that the critical behavior of correlation functions is affected significantly, reflecting a scenario that is usually encountered in physical systems exhibiting fractionalization. The latter is a property underlying the decay of an order parameter into emergent building blocks, similarly to the physics of quark deconfinement as studied in high-energy physics. In condensed matter physics, it is typically associated to an emerging gauge symmetry, with realizations in quantum antiferromagnets featuring competing orders. This typically leads to a phase structure that is incompatible with the LGW paradigm [1, 4]. The corresponding critical behavior where competing order parameters decay into fractionalized quasi-particles is what one generally refers to DQC. One of the most salient features underlying this scenario is that it leads to a unusually large anomalous dimension, $\eta$, of the order parameter $O(\bm{x})$ [1]. Recall that in LGW-compatible systems $\eta$ is typically considerably smaller than the thermodynamic critical exponents, namely, those associated to the critical singularities of the free energy density.
In this Letter, we show that accounting for the constraint ${\mbox{$\nabla$}}\cdot\bm{P}=0$ implies a large anomalous dimension, with no obvious presence of fractionalized excitations. In fact, we find at two-loop order in perturbation theory that $\eta=0.239$ in three dimensions. By contrast, in a LGW theory with $O(N)$ symmetry one finds a value of $\eta$ which is typically one order of magnitude smaller than the value found here. Indeed, in an $O(3)$ symmetric theory, which is the natural case to compare with in the present context, the result is $\eta\approx 0.034$. Crucially, the constraint ${\mbox{$\nabla$}}\cdot\bm{P}=0$ naturally imposes that $N=D$, i.e., the number of components of the order parameter coincides with spatial dimensionality.
@@
Substituting the fixed point values, $\hat{a}_{*}=-25/77$ and $\hat{b}_{*}=5/17$, calculated from Eqs. (9) for $D=3$, we obtain $\eta\approx 0.239$. This value is significantly larger than the ones typically obtained from the LGW theory, thus highlighting the impact of the constraint ${\mbox{$\nabla$}}\cdot\bm{P}=0$.
@@
It remains to be seen whether the universality class uncovered here can be connected to some fractionalization mechanism, by representing $\bm{P}$ as a composite operator, similarly to the CP1 representation in quantum antiferromagnets [1]. DQC is indeed facilitated by a mechanism where staggered Berry phases interfere with quadrupled instanton events, which ultimately causes the order parameter to fall apart into spinons [1, 4].
Conclusions and final remarks – In this Letter, we have shown that imposing a local divergence-free constraint on the polarization field drives a ferroelectric system into a universality class that lies beyond the conventional LGW framework. This profoundly modifies the configuration space, restricting it to loop-like polarization structures and thereby altering the critical behavior. The RG analysis demonstrates that the polarization correlation function features a large anomalous dimension in three dimensions, $\eta=0.239$, which significantly exceeds that of ordinary $O(3)$ LGW models.
This unconventional scaling originates from an emergent $U(1)$ gauge symmetry, which becomes explicit when the constraint is resolved as $\bm{P}={\mbox{$\nabla$}}\times\bm{A}$, rendering the effective theory manifestly gauge invariant. In this sense, the zero divergence constraint is understood as a local conservation law corresponding to a local gauge symmetry. This leads to the emergence of a gauge field and therefore to a large value of $\eta$. Usually, the argument runs the opposite way, with a gauge symmetry yielding a local conservation law, one example being current conservation in electrodynamics. From a field theoretic perspective, at the critical point current conservation implies that currents do not have an anomalous dimension [40]. This fact is a consequence of the scale invariance of the critical point. The opposite is true here: the order parameter, viewed as a conserved quantity, has a large anomalous dimension.
Looking ahead, ferroelectric systems that naturally enforce local polarization constraints [14] open a route for exploring gauge-symmetry mediated universality and constraint-induced critical behavior in classical systems.